# Fine-Tuning a Time Series Foundation Model with Wasserstein Loss

### v1

@@ 3 Wasserstein Deep Learning Model for Tokenized Time Series
In this section, we discuss our approach to applying Wasserstein loss to a tokenized deep learning model. We believe that our method can be applied to any task where the distance between classes plays a significant role.
@@ 3.3 Loss Function
$W_{p}(Y_{a},\hat{Y})=r\cdot\left(\sum_{i=1}^{d}p_{i}\cdot|i-a|^{p}\right)^{1/p}$ (2)
where $Y_{a}$ represents the ground truth degenerate distribution, equal to the token $a$, $\hat{Y}$ is the forecasted distribution over tokens, $d$ is the number of tokens, and $p_{i}$ is the predicted probability for token $i$. Note that if the model output were a scalar, this formula would simplify to well-known regression losses: mean absolute error when $p=1$ and root mean squared error when $p=2$.
@@ 4.2 Fine-Tuning Results
As discussed in Section 3, we fine-tuned the pretrained Chronos-T5 (Small) model. For each dataset, we conducted $1000$ fine-tuning steps, with the initial learning rate set to $0.001$, which linearly decreased to 00 over the course of the steps. We fine-tuned the model using three different loss functions. The first two, Wasserstein-1 (W1) and Wasserstein-2 (W2), correspond to equation 2, with $p=1$ and $p=2$, respectively. The third loss function is the standard cross-entropy loss. Additionally, we calculated metrics for the model without fine-tuning.
Figures 2 and 2 present the results for MASE and WQL, respectively. Appendix A provides the MASE and WQL values for each dataset. The Wasserstein loss significantly outperforms cross-entropy loss in point estimation; however, we observe a degradation in the WQL metric. This is a direct result of the loss design. Since we use a degenerate distribution as the target in the Wasserstein loss, the forecasted distribution becomes sharper and less suitable for quantile estimation compared to cross-entropy loss.
@@ 5 Discussion
The primary follow-up to this paper would be to train a foundational time series model from scratch using Wasserstein loss. Although probabilistic forecasting was not the main focus of this work, the model’s ability to capture uncertainty is crucial. Improvements in probabilistic forecasting could be achieved by relaxing the assumption that the target distribution is a degenerate distribution.
Appendix A Appendix

### v2

@@ 3 Wasserstein Deep Learning Model for Tokenized Time Series
In this section, we discuss our approach to applying Wasserstein loss to a tokenized deep learning model. Our method can be generalized to tasks where the relative distance between predicted classes is crucial.
@@ 3.3 Loss Function
$W_{p}(Y_{a},\hat{Y})=r\cdot\left(\sum_{i=1}^{d}\alpha_{i}\cdot|i-a|^{p}\right)^{1/p}$ (2)
where $Y_{a}$ represents the ground truth degenerate distribution, equal to the token $a$, $\hat{Y}$ is the forecasted distribution over tokens, $d$ is the number of tokens, and $\alpha_{i}$ is the predicted probability for token $i$. Note that if the model output were a scalar, this formula would reduce to well-known regression losses: the absolute error (AE) when $p=1$ and the squared error (SE) when $p=2$. The relationship between Wasserstein losses and AE/SE is explored further in Appendix B.
@@ 4.2 Fine-Tuning Results
As discussed in Section 3, we fine-tuned the pretrained Chronos-T5 (Small) model [^2] The code is available at https://github.com/ChernovAndrey/chronos-forecasting-wasserstein.git. For each dataset, we conducted $1000$ fine-tuning steps, with the initial learning rate set to $0.001$, which linearly decreased to 00 over the course of the steps. We fine-tuned the model using three different loss functions. The first two, Wasserstein-1 (W1) and Wasserstein-2 (W2), correspond to equation 2, with $p=1$ and $p=2$, respectively. The third loss function is the standard cross-entropy loss. Additionally, we calculated metrics for the model without fine-tuning.
Figures 2 and 2 present the results for MASE and WQL, respectively. Appendix A provides the MASE and WQL values for each dataset. The Wasserstein loss significantly outperforms cross-entropy loss in point estimation; however, we observe a degradation in the WQL metric. This is a direct result of the loss design. Since we use a degenerate distribution as the target in the Wasserstein loss, the forecasted distribution becomes sharper and less suitable for quantile estimation compared to cross-entropy loss. See Appendix C for futher details.
@@ 5 Discussion
A key area for future work is training a foundational time series model from scratch using Wasserstein loss. Although probabilistic forecasting was not the main focus of this work, the model’s ability to capture uncertainty is crucial and should be explored in future research.
Appendix A Results for each dataset
@@
Appendix B Relationship between Wasserstein Loss and Common Regression Losses
In this section, we demonstrate that the absolute error (AE) and squared error (SE) of a forecast’s expected value and its target, defined as follows:
$\text{AE}(Y_{a},\mathbb{E}[\hat{Y}])=|\mathbb{E}[\hat{Y}]-a|,\quad\text{and}\quad\text{SE}(Y_{a},\mathbb{E}[\hat{Y}])=(\mathbb{E}[\hat{Y}]-a)^{2},$
serve as lower bounds for the Wasserstein losses:
$W_{p}(Y_{a},\hat{Y})=r\cdot\left(\sum_{i=1}^{d}\alpha_{i}\cdot|i-a|^{p}\right)^{1/p}=r\cdot\left(\mathbb{E}[|i-a|^{p}]\right)^{1/p}\geq r\cdot\left(|\mathbb{E}[Y]-a|^{p}\right)^{1/p},$ (3)
where $p\in\{1,2\}$. The last inequality follows from Jensen’s inequality, since the function $f(i)=|i-a|^{p}$ is convex for $p=1$ or $p=2$.
Thus, we obtain:
$W_{1}(Y_{a},\hat{Y})\geq\text{AE}(Y_{a},\mathbb{E}[\hat{Y}])\quad\text{and}\quad W_{2}(Y_{a},\hat{Y})^{2}\geq\text{SE}(Y_{a},\mathbb{E}[\hat{Y}]).$
Although we did not conduct experiments using AE and SE losses, exploring these could be a promising direction for future research.
Appendix C Distribution Forecasting
Figure 3: Kernel density estimation (KDE) comparison of forecasts on the FRED-MD dataset. The plot shows that the model trained with W1 loss produces significantly sharper forecast distributions compared to the model trained with cross-entropy loss.
In this section, we explore why models trained with the Wasserstein loss tend to exhibit worse WQL metrics compared to those trained with cross-entropy loss.
The asymptotic behavior of these loss functions diverges significantly as the number of tokens ($n$) increases. Consider a simple case where the predicted distribution consists of only two tokens with non-zero probabilities: $p$ and $1-p$, where $p\in(0,1)$. Additionally, assume that the non-zero tokens are the first and last in the sequence, with the target corresponding to the token with probability $p$. In this scenario, the cross-entropy loss remains constant as $n$ increases, given by $-\log(p)$. In contrast, the Wasserstein loss depends on $n$ and diverges as $n$ approaches infinity, following $((1-p)\cdot(n-1)^{p})^{1/p}$. This behavior resembles that of regression losses, resulting in a sharper predicted distribution, as illustrated in Figure 3. Consequently, when the model’s predictions deviate significantly from the target, the quantile loss increases substantially.
We investigated these differences in forecasting using the Monash FRED-MD dataset, which contains 107 monthly time series representing various macroeconomic indicators from the Federal Reserve Bank. Our analysis showed that the model trained with the W1 loss performed better on 55 time series, while the model trained with the cross-entropy (CE) loss performed better on 52 time series. However, the aggregated WQL metric of the model trained with CE loss was significantly better. This is because, when the model’s predictions deviate notably from the ground truth, the quantile losses for the model trained with W1 loss are much higher. Nonetheless, since models trained with Wasserstein loss exhibit less bias in point estimation, incorporating sampling techniques like Monte Carlo Dropout could potentially enhance the quality of distribution forecasts.